OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 9x + y = 0 18x + 4y = -8

OpenStudy (anonymous):

how would you solve it? explain please

OpenStudy (anonymous):

rewrite the first one as \[y=-9x\] and then replace \(y\) in the second equation with \(-9x\) and solve \[18x+4\times (-9x)=-8\] for \(x\)

OpenStudy (anonymous):

18x+-36x=-8

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

x=4/9? @satellite73

OpenStudy (anonymous):

i don't know, i didn't do it \[-18x=-8\] \[x=\frac{8}{18}=\frac{4}{9}\] yeah looks good

OpenStudy (anonymous):

how do i solve for y @satellite73

OpenStudy (anonymous):

i will let you do that on your own hint: \[9x+y=0\] and \[x=\frac{4}{9}\]

OpenStudy (anonymous):

plug in the x , but then what @satellite73